# A Holistic Optimization Framework for Energy Efficient UAV-assisted Fog Computing: Attitude Control, Trajectory Planning and Task Assignment

### v2

@@ VI-B1 Optimal Propeller Parameters
Figure 7: Relationship between overall energy consumption and propeller blades parameters.
VI-B2 Convergence Speed of Algorithms
Fig. 8 demonstrates the average convergence speeds of $200$ independent runs by the classical ACS, ACS-D and ACS-DS over $80$ iterations. Both mechanisms, decoupling and safety values, reduce the energy consumption of UAV movement to an extent. In addition, with classical ACS as the baseline, ACS-D improves the convergence speed to obtain the optimal trajectory by more than $7\%$, while the improvement by ACS-DS is more than $15\%$.
@@ VI-B2 Convergence Speed of Algorithms
VI-B3 Trajectory planning
Figure 9: A demonstration of trajectories planned by different mechanisms on a 3D spatial domain with terrain height generated randomly as a continuous surface.
We now generate the terrain height randomly as a continuous surface (no-fly zone) within the three-dimensional spatial domain as shown in Figs. 9 and 10, and consider that the MDs are placed directly on the ground conforming to the terrain height, in order to test the obstacle avoidance capability of our proposed trajectory planning and attitude control algorithms. We observe from the trajectories of the UAV in Figs. 9 and 10 that, both the additional mechanisms of decoupling and safety values in ACS and attitude control (ATC) have positive impacts on the final path. Specifically, ATC achieves more stable and precise flight control by optimizing the UAV rotor speed in real time based on the feedback of error and environmental information from the sensors. By jointly applying the ACS with the decoupling, safety values and attitude control, ACS-DS+ATC obtains the smoothest and shortest trajectory among all.
Figure 10: Demonstration of trajectories planned by different mechanisms on a 3D spatial domain with terrain height generated randomly as a continuous surface (TOP).
VI-B4 Operational Efficiency Cost
Finally, we present the results of the operational efficiency cost (combining delay and energy consumption) achieved by the eight implementations in Fig. 11. Again, we can observe that the decoupling mechanism and safety values in ACS-DS can significantly reduce the overall consumption compared to the classical ACS.
In addition, our proposed attitude control mechanism (ATC) can further improve the performance, irrespective of whether ACS, ACS-D, or ACS-DS is employed. Overall, ACS-DS+ATC can reduce the operational efficiency cost by more than $58\%$ and $35\%$ compared to RAN and classical ACS, respectively.
Figure 11: Comparison of operational efficiency cost by different methods.
VII Conclusions
@@
In this paper, we have proposed a joint optimization framework to reduce the operational efficiency cost in UAV-assisted fog computing systems. Our proposed framework involves multiple modules, including quadrotor UAV attitude control, trajectory planning in a three-dimensional spatial domain with continuously varying terrain heights, and energy-efficient assignment of computing tasks to different components in the network. We have designed appropriate mechanisms or algorithms for each module in the framework, and integrated them together to obtain a holistic solution to improve the overall efficiency in UAV-assisted fog computing. Specifically, we have proposed a novel fuzzy PID control mechanism for effective attitude control, designed the ACS-DS algorithm that overcomes the convergence issue in conventional approaches for trajectory planning in three-dimensional domains, and a modified PSO algorithm to determine the optimal task assignment. Numerical results from a wide range of experiments have shown that our proposed framework can reduce the operational efficiency cost by over $35\%$ compared to existing approaches.
-A The Dynamics of a Quadrotor UAV
@@
  • The stator magnetic field speed and rotor speed of the motors are infinitely close to each other, such that the slip rate is 00;
@@
  • The Euler angles are bounded, i.e., $-\pi/2<\phi<\pi/2,-\pi/2<\theta<\pi/2,-\pi<\psi<\pi$.
@@
$\begin{array}[]{l}R_{b}^{e}=\left(R_{e}^{b}\right)^{-1}=\left(R_{e}^{b}\right)^{T}=\\ \\ \left[\begin{array}[]{cll}\cos\theta\cos\psi&\sin\theta\cos\psi\sin\phi&\sin\theta\cos\psi\sin\phi\\ &-\sin\psi\cos\phi&+\sin\psi\cos\phi\\ \cos\theta\sin\psi&\sin\theta\sin\psi\sin\phi&\sin\theta\sin\psi\sin\phi\\ -\sin\theta&+\cos\psi\cos\phi&-\cos\psi\cos\phi\\ &\cos\theta\sin\phi&\cos\theta\cos\phi\end{array}\right]\end{array}$ (27)
Four inputs, including pitch control quantity $U_{1}$, roll control quantity $U_{2}$, yaw control quantity $U_{3}$, and altitude control quantity $U_{4}$, can be controlled based on the torque of the four motors of the UAV, that is,
$\left\{\begin{array}[]{l}U_{1}(t)=C_{T}(t)\left(\omega_{1}(t)^{2}+\omega_{2}(t)^{2}+\omega_{3}(t)^{2}+\omega_{4}(t)^{2}\right)\\ U_{2}(t)=C_{T}(t)\left(-\omega_{2}(t)^{2}+\omega_{4}(t)^{2}\right)\\ U_{3}(t)=C_{T}(t)\left(-\omega_{1}(t)^{2}+\omega_{3}(t)^{2}\right)\\ U_{4}(t)=C_{M}(t)\left(-\omega_{1}(t)^{2}+\omega_{2}(t)^{2}-\omega_{3}(t)^{2}+\omega_{4}(t)^{2}\right)\end{array}\right.$ (28)

### v3

@@ 6.4.1 Optimal Propeller Parameters
Figure 9: Overall energy consumption vs. propeller blades parameters.
6.4.2 Convergence Performance
Fig. 10 demonstrates the average convergence speeds of GA-SCA, CPS-ACO, and ACS-DS over $60$ iterations, and TD3 algorithm over $3000$ episodes for $200$ independent runs. The results show that both ACS-DS and TD3 can optimize the energy efficiency through efficient exploring optimal trajectories. However, the average training time per episode for the RL-based TD3 are significantly higher than the average running time per iteration of the other three heuristic algorithms. Therefore, due to its faster convergence and lower computational overhead, ACS-DS is more suitable for rapid decision-making and trajectory planning in unknown environments.
@@ 6.4.2 Convergence Performance
6.4.3 Trajectory planning with attitude control
From the trajectories shown in Figs. 7(a) and 7(b), we observe that, the CPS-ACO, TD3, and ACS-DS algorithms are more capable of avoiding no-fly zones. In addition, our proposed Attitude Control (ATC) method can improve the trajectories planned by all algorithms, by effectively correcting unwanted directional changes. Specifically, the FEAR-PID controller dynamically optimizes flight attitude by adjusting the UAV rotor speed in real-time based on sensor feedback, including error and environmental data, resulting in more stable and precise flight control.
6.4.4 Operational Efficiency Cost
We present the results of the operational efficiency cost (considering both delay and energy consumption) achieved by different implementations in Fig. 11. Note that the horizontal axis in this figure (and similar plots throughout this section) represents discrete control time slots, each of duration $L=0.1$ s as defined in Section 3, rather than the UAV’s actual physical flight time. The overall mission duration $T$ spans 20 to 35 minutes in our experiments, calculated as described in Section 6.1 below Table 5), and the figures display representative segments of the trajectory for clarity of visualization. One important observation is that, the decoupling mechanism and safety values in ACS-DS can significantly reduce the overall consumption compared to the classical ACS, and have a slight advantage over SOTA (state-of-the-art) reinforcement learning methods such as TD3.
Moreover, the ATC mechanism further enhances the performance of both RL-based (TD3) and heuristic-based (ACS-DS and CPS-ACO) methods. Overall, ACS-DS+ATC achieves the best performance in terms of operational efficiency, closely followed by TD3+ATC. Specifically, compared with conventional ACS, TD3+ATC reduces the operational efficiency cost by 43.5%, while ACS-DS+ATC reduces it by 48.1%.
These results validate our initial claim in this paper, that jointly optimizing attitude control, trajectory planning, and task assignment in UAV-assisted fog computing systems effectively captures the interdependencies among these modules. Such integrated approaches (e.g., TD3+ATC and our proposed ACS-DS+ATC) lead to significant improvements in system performance compared to optimizing each aspect individually.
Figure 11: Comparison of operational efficiency cost.
6.4.5 Trajectory Stability and Robustness Evaluation
To evaluate the stability and robustness of different control and planning algorithms in complex 3D environments, we conducted $50$ independent runs for each method under identical initial conditions and environmental settings. The trajectories were generated by our Framework, validated in Isaac Sim and AirSim, and subsequently rendered in the WebGL visualization environment to provide a clearer and lightweight presentation. In all figures, each polyline corresponds to one complete trajectory obtained from a single run.
Fig. 12 compares the tracking performance of the classical PID controller (red) and the proposed FEAR-PID controller (blue) along the same reference path. The trajectories generated by PID exhibit noticeable drift and oscillation, especially in densely cluttered urban regions. In contrast, FEAR-PID produces a significantly more compact trajectory cluster with reduced lateral deviation, demonstrating superior attitude stability and higher run-to-run consistency.
Fig. 13 presents the trajectory distributions of three planning algorithms: ACS (yellow), TD3 (green), and ACS-DS (blue). ACS exhibits the largest dispersion, indicating higher sensitivity to environmental variations. TD3 shows moderate improvement in stability, while ACS-DS produces the most concentrated trajectories among the three, confirming that its decoupling mechanism and dynamic safety factors substantially enhance planning robustness across repeated trials.
Overall, these results show that,
  1. 1.
    FEAR-PID effectively reduces disturbance-induced drift and improves attitude control precision;
  2. 2.
    ACS-DS generates more stable and reliable flight paths under repeated execution;
  3. 3.
    Jointly optimizing control and planning is essential for ensuring dependable UAV operation in realistic 3D environments.
Figure 12: Trajectory comparison between PID (red) and FEAR-PID (blue). Visualization of $50$ independent runs for evaluating trajectory stability and robustness.
Figure 13: Trajectory comparison of ACS (yellow), TD3 (green), and ACS-DS (blue). Visualization of $50$ independent runs for evaluating trajectory stability and robustness.
6.4.6 Ablation Study Results
We summarize the ablation results in Fig. 14, which report the average operational efficiency cost and its variance across configurations. The full model consistently achieves the lowest cost. Removing the ACS-DS trajectory planner causes the sharpest degradation, reflecting its importance in balancing energy and delay. The absence of PSO-based task assignment also leads to significant overhead due to poor task-resource coordination.
Interestingly, even the removal of the FEAR-PID attitude controller, while seemingly less impactful than trajectory and assignment modules, causes noticeable inefficiencies, particularly in maneuvering-intensive segments such as ascent, descent, and turns. This confirms that flight stability indirectly influences task execution and communication quality. The gamma-based channel allocation strategy, though relatively lightweight, also contributes to system-wide efficiency by avoiding bandwidth contention and ensuring smoother transmission.
Overall, the results confirm that all modules are complementary. The holistic approach offers not only the best mean performance but also the most stable behavior, reinforcing the need for joint optimization in UAV-assisted fog computing.
Figure 14: Average operational efficiency cost across ablation variants.
7 Conclusions
@@
In this paper, we have proposed a joint optimization framework to reduce the operational efficiency cost in UAV-assisted fog computing systems. Our proposed framework involves multiple modules, including quadrotor UAV attitude control, trajectory planning in a three-dimensional spatial domain with continuously varying terrain heights, and energy-efficient assignment of computing tasks to different components in the network. We have designed appropriate mechanisms or algorithms for each module in the framework, and integrated them together to obtain a holistic solution to improve the overall efficiency in UAV-assisted fog computing. Specifically, we have proposed a novel FEAR-PID control mechanism for effective attitude control, designed the ACS-DS algorithm that overcomes the convergence issue in conventional approaches for trajectory planning in three-dimensional domains, and a modified PSO algorithm to determine the optimal task assignment. Numerical results from a wide range of experiments have shown that our proposed framework can reduce the operational efficiency cost significantly compared to existing approaches.
Appendix A The Dynamics of a Quadrotor UAV
@@
  • 3.
    The stator magnetic field speed and rotor speed of the motors are infinitely close to each other, such that the slip rate is 0;
  • 4.
@@
  • 5.
    The Euler angles are bounded, i.e., $\pi/2<\phi<\pi/2,-\pi/2<\theta<\pi/2,-\pi<\psi<\pi$.
@@
$\begin{aligned} R_{b}^{e}&=\left(R_{e}^{b}\right)^{-1}=\left(R_{e}^{b}\right)^{T}=\\ &\begin{bmatrix}\cos\theta\cos\psi&\cos\psi\sin\theta\sin\phi-\sin\psi\cos\phi&\cos\psi\sin\theta\cos\phi+\sin\psi\sin\phi\\ \cos\theta\sin\psi&\sin\psi\sin\theta\sin\phi+\cos\psi\cos\phi&\sin\psi\sin\theta\cos\phi-\cos\psi\sin\phi\\ -\sin\theta&\sin\phi\cos\theta&\cos\phi\cos\theta\end{bmatrix}\end{aligned}$
Four inputs, including total thrust $U_{1}$, roll torque $U_{2}$, pitch torque $U_{3}$, and yaw torque $U_{4}$, can be controlled based on the rotation speeds of the four motors of the UAV, that is,
$\left\{\begin{array}[]{l}U_{1}(t)=C_{T}(t)\left(\omega_{1}(t)^{2}+\omega_{2}(t)^{2}+\omega_{3}(t)^{2}+\omega_{4}(t)^{2}\right)\\ U_{2}(t)=C_{T}(t)\left(-\omega_{2}(t)^{2}+\omega_{4}(t)^{2}\right)\\ U_{3}(t)=C_{T}(t)\left(-\omega_{1}(t)^{2}+\omega_{3}(t)^{2}\right)\\ U_{4}(t)=C_{M}(t)\left(-\omega_{1}(t)^{2}+\omega_{2}(t)^{2}-\omega_{3}(t)^{2}+\omega_{4}(t)^{2}\right)\end{array}\right$.